OpenStudy (anonymous):

Evaluate the integral Integral of (x^3)/(1+x^2)^(1/2). I thought you would use x=tan of theta, but that doesn't seem to work, unless I messed up somewhere.

OpenStudy (anonymous):

1/3 (x^2-2)sqrt(X^2+1)

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

Ya, I know what the answer is. I just don't know how to get to it.

OpenStudy (anonymous):

you can use the u subsitution

OpenStudy (anonymous):

u=x^2+1 du=2x dx du/2=dx

OpenStudy (anonymous):

or u=x^2 du=2x dx 1/2 integral u/sqrt(u+1)

OpenStudy (anonymous):

there are many ways to do this you can also rewrite thequestion as x^2*x/sqrt(x^2+1)

OpenStudy (anonymous):

Would using x=tan of theta work?

OpenStudy (jamesj):

It would, but it would just make the problem more complicated. Why use a trig substitution when you can just use a polynominal substitution?